If L is a line perpendicular to 2y-3x=1, then the sum of the slopes of the two lines is:
A) 5/6
B)3/2
c)2/3
D)-1
E)1/5

Answer by fractalier(6550)

The slope of 2y-3x=1 can be found by solving for y...
2y = 3x + 1
y = (3/2)x + 1/2
Its slope is 3/2.
Its perpendicular slope is -2/3.
Adding the we get
3/2 - 2/3 = 5/6
Choice A.